# Electronic order induced symmetry breaking in lattice dynamics of Co3Sn2S2

### v2

@@ Abstract
MBC-modified $E_{u}$ mode Figure 4(a) displays the MBC-modified phonon spectra of the $E_{u}$ mode along the path $\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$. The MBC also lifts the degeneracy of the $E_{u}$ mode (dashed line) across the entire $\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$ path. The magnitude of the phonon energy splitting for the Eg modes is presented in Fig. 4(b), revealing a maximum splitting in the vicinity of T and minimum splitting in the vicinity of $\Gamma$.
Combined with Figs. 3(b) and  4(b), it can be observed that although the parity of the $E_{g}$ and $E_{u}$ modes remains unchanged at both $\Gamma$ and T, the relative phase between the two S atoms shifts by $\pi$. For instance, the $E_{g}$ mode exhibits a relative phase of $\pi$ at $\Gamma$ but 0 at T, whereas the $E_{u}$ mode shows a phase of 0 at $\Gamma$ and $\pi$ at T. This switching of the relative phase between the two modes correlates with the reversal of the maximum phonon splitting.
@@
At the $\Gamma$ point, the splitting calculated from our algorithm is about 0.002 meV (0.016 cm-1), which is smaller than the experimentally observed value of 0.3 cm-1 reported in the optical conductivity measurements [9]. This discrepancy suggests that the splitting observed in experiments is not primarily caused by the MBC. Combined with the asymmetric line shapes of the $E_{u}$ mode, we attribute the larger experimental splitting to a Fano resonance [9]. This resonance may arise from quantum interference between electronic and phonon excitations when Weyl cones approach the Fermi level [45, 46]. The absence of an asymmetric line shape in the $E_{g}$ mode further supports this interpretation, indicating that the phonon splitting mechanisms for the $E_{g}$ and $E_{u}$ modes are distinct.
We note that the energy splitting of the $E_{u}$ mode is significantly smaller than that of the $E_{g}$ mode at $\Gamma$ point, a result captured by both our ab initio calculations and experimental measurements [8, 9]. Moreover, We notice both the $E_{g}$ and $E_{u}$ modes exhibit a broad distribution of splitting across the entire BZ, rather than being concentrated solely at the high-symmetry momentum. This widely distributed chiral phonon dynamics may give rise to observable transport signatures associated with phonon chirality.
@@
Fig. 4(c) displays the distributions of pseudo-angular momentum and angular momentum for the two split $E_{u}$ modes. The angular momentum of these modes is not quantized to unity, which can be attributed to the vibration from both Co and S atoms. As illustrated in Fig. 4(d), the Co atoms are not located at $C_{3}$-invariant Wyckoff positions, thus resulting in a non-quantized values of angular momentum [23].
@@
Figure 5: Distribution of the MBC in the electronic BZ. (a-c) are MBC distributions for the $E_{g}$ phonon mode with $l_{ph}=+1$ on three distinct $k_{z}$ planes. The MBC distribution for the $E_{g}$ mode with $l_{ph}=-1$ is of opposite sign. (d-f) are MBC distributions for the $E_{u}$ mode with $l_{ph}=+1$. The $E_{u}$ mode with $l_{ph}=-1$ also exhibits a sign reversal. The MBC magnitude is significantly larger for the $E_{g}$ modes than for the $E_{u}$ modes, which consequently results in a smaller calculated phonon splitting for the $E_{u}$ modes.
@@
Phonon splitting and magnetism under MBC To explain the distinct magnitude of the energy splitting for the $E_{g}$ and $E_{u}$ modes at the $\Gamma$ point, we calculate the distribution of the mode-resolved MBC, denoted as $\tilde{G}(\Gamma,\bm{k},\bm{k})$, across the electronic BZ, as shown in Fig. 5. The results indicate that $\tilde{G}_{E_{u}}(\Gamma,\bm{k},\bm{k})$ is significantly smaller in magnitude than $\tilde{G}_{E_{g}}(\Gamma,\bm{k},\bm{k})$, consistent with our earlier findings on MBC-induced phonon splitting. Furthermore, we observe that the distribution of $\tilde{G}_{E_{u}}(\Gamma,\bm{k},\bm{k})$ exhibits substantial cancellation between positive and negative contributions. This cancellation further explains the markedly smaller MBC-induced splitting of the $E_{u}$ mode compared to the $E_{g}$ mode.
@@
Previous studies have attributed $\Gamma$-point phonon splitting under external magnetic fields to the manifestation of an effective phonon magnetic moment ($M_{z}$[3, 4, 1, 2]. As detailed in Ref. [3], $M_{z}$ comprises two contributions: one from the Born effective charges and another from the Berry curvature of electrons in both real space (MBC) and momentum space ($\Omega_{e}$), i.e.,
$M_{z}=\frac{e}{2M_{\kappa}}l_{z,\nu\bm{q}}\int d\bm{k}\left\{\Omega_{k_{x}u_{y}}-\Omega_{k_{y}u_{x}}+\Omega_{e,k_{x}k_{y}}G_{u_{x}u_{y}}\right\}.$ (4)
@@
The first two terms collectively integrate to the Born effective charge, whereas the third term is equivalent to the product of the conventional electronic and molecular Berry curvatures. In css, the metallic nature and the preservation of $\mathcal{P}$ result in a vanishing Born effective charge. Given that $G$ is independent of $\bm{k}$ and the integral $\int d\bm{k}\Omega_{e,k_{x}k_{y}}$ is proportional to the anomalous Hall conductivity $\sigma^{\text{A}}_{xy}$, the expression for the phonon magnetic moment simplifies to $M_{z}\propto l_{z,\nu\bm{q}}\cdot G\cdot\sigma^{\text{A}}_{xy}$. Consequently, the large intrinsic Hall conductivity of css (approximately 505 $\Omega^{-1}$ cm-1 [25]) indicates a strong propensity for intrinsic phonon magnetism in this material. As calculated and shown in Fig. 5, the pronounced difference in the MBC between the $E_{g}$ and $E_{u}$ modes indicates a larger magnetic moment for the $E_{g}$ phonon.
@@
In summary, our ab initio-MBC methodology captures the electronic order-induced symmetry breaking in the lattice dynamics of css, generalizing the manifestation of phonon magnetism and phonon energy shift effects in magnetic materials. Notably, the energy scale of circularly polarized phonons coincides with that of chiral Weyl fermions, suggesting rich opportunities for chiral boson–fermion coupling and enabling control over topological states through phononic and magnetic means [47]. This efficient and accurate framework for magnetic lattice dynamics is readily extendable to full Brillouin-zone calculations and enables the study of emergent phenomena such as the phonon Hall effect [13, 14, 15, 16, 17, 15, 18, 19, 20, 4], Hall viscosity [48], and so on. The algorithm is broadly applicable to intrinsic magnetic systems, facilitating the discovery of materials with large phonon magnetism.
Our findings establish a set of design principles for materials with large phonon magnetic moments and substantial splitting, i.e., strong spin-orbit coupling, significant electron-phonon coupling, and phonon modes carrying nonzero angular momentum. We also propose that materials integrating magnetic order with flat bands, such as css, represent a particularly fertile platform for the discovery of giant phonon splitting and pronounced phonon magnetism.
Future studies may focus on extrinsic contributions to the phonon magnetic moment, such as relaxation processes [4], non-reciprocal transport related to circularly polarized phonons [49, 21, 50, 51], and ab initio algorithm for external field–induced phonon magnetism, etc.
I Acknowledgement
@@
We acknowledge the helpful discussion with Haijun Zhang, Fuyi Wang, Xi Dai, Yang Gao, Yafei Ren, Shang Ren, Run Yang and Martin Dressel. We acknowledge the support from National Key R&D Project (Grant Nos. 2023YFA1407400 and 2024YFA1409200), and the National Natural Science Foundation of China Grant No. 12374165.

### v3

@@ II.2.2 MBC-modified $E_{u}$ mode
Figure 2(e1) displays the MBC-modified phonon spectra of the $E_{u}$ mode along the path $\text{T}\leftrightarrow\Gamma\leftrightarrow\text{T}$, where the MBC also lifts the degeneracy (dashed line) across the entire path. The magnitude of the phonon energy splitting for the $E_{u}$ modes is presented in Fig. 2(e2), revealing a maximum splitting in the vicinity of T and minimum splitting in the vicinity of $\Gamma$. Both $E_{g/u}$ modes exhibit a broad distribution of splitting across the entire BZ, rather than being concentrated solely at the high-symmetry momentum. This widely distributed chiral phonon dynamics may give rise to observable transport signatures associated with phonon chirality.
@@
At the $\Gamma$ point, the splitting calculated from our algorithm is about 0.002 meV (0.016 cm-1), which is smaller than the experimentally observed value of 0.3 cm-1 reported in the optical conductivity measurements [CoSnS_CP_2025]. This discrepancy suggests that the splitting observed in experiments is not primarily caused by the MBC. Combined with the asymmetric line shapes in experiments, we attribute the larger experimental splitting to the Fano resonance, which may arise from quantum interference between electronic and phonon excitations when Weyl cones approach the Fermi level [cappelluti_charged2012, CoSnS_CP_2025]. The absence of an asymmetric line shape in the $E_{g}$ mode further supports this interpretation, indicating that the phonon splitting mechanisms for the $E_{g}$ and $E_{u}$ modes are distinct.
@@
Figure 2(e3) displays the distributions of $l_{ph}$ and $l_{z,\nu\bm{q}}$ for the two split $E_{u}$ modes, where $l_{z,\nu\bm{q}}$ is not quantized to unity. This can be attributed to the vibration from both Co and S atoms, and Co atoms are not located at $C_{3}$-invariant Wyckoff positions [Zhang2025_CP], as illustrated in Fig. 2(e4).
@@
Figure 3: Distribution of the MBC in the electronic BZ. (a-c) are MBC distributions for the $E_{g}$ phonon mode with $l_{ph}=+1$ on three distinct $k_{z}$ planes. The MBC distribution for the $E_{g}$ mode with $l_{ph}=-1$ is of opposite sign. (d-f) are MBC distributions for the $E_{u}$ mode with $l_{ph}=+1$. The $E_{u}$ mode with $l_{ph}=-1$ also exhibits a sign reversal. The MBC magnitude is significantly smaller for the $E_{u}$ modes, which consequently results in a smaller calculated phonon splitting.
II.3 Phonon splitting and magnetism from MBC
@@
For css, the energy splitting of the $E_{u}$ mode is significantly smaller than that of the $E_{g}$ mode at $\Gamma$ point, a result captured by both our ab initio calculations and experimental measurements [CoSnS_PRL_2025, CoSnS_CP_2025]. To explain their distinct magnitude of the energy splitting, we calculate the distribution of the mode-resolved MBC, denoted as $\tilde{G}(\Gamma,\bm{k},\bm{k})$, across the electronic BZ, as shown in Fig. 3. The results indicate that $\tilde{G}_{E_{u}}(\Gamma,\bm{k},\bm{k})$ is significantly smaller in magnitude than $\tilde{G}_{E_{g}}(\Gamma,\bm{k},\bm{k})$, consistent with our earlier findings on MBC-induced phonon splitting. Furthermore, we observe that the distribution of $\tilde{G}_{E_{u}}(\Gamma,\bm{k},\bm{k})$ exhibits substantial cancellation between positive and negative contributions, which further explains the markedly smaller MBC-induced splitting of the $E_{u}$ mode.
We note that although the $E_{u}$ modes involves vibrations from both Co and S atoms, the atom-resolved MBC shows that S atom vibrations dominate its contribution (details are in the Supplementary Material Fig. LABEL:fig:Eu_seperate). Combined with Figs. 2(d2) and  2(e2), it can be observed that while the parity of the $E_{g/u}$ modes remains unchanged at both $\Gamma$ and T, the relative phase between the two S atoms shifts by $\pi$. For instance, the $E_{g}$ mode exhibits a relative phase of $\pi$ at $\Gamma$ but 0 at T, whereas the $E_{u}$ mode shows a phase of 0 at $\Gamma$ and $\pi$ at T. The distribution of the maximum phonon splitting for the $E_{g/u}$ modes switches along the $\Gamma$–T path, a reversal that correlates with the relative phase between two S atoms of the modes.
@@
Previous studies have attributed $\Gamma$-point phonon splitting under external magnetic fields to the manifestation of an effective phonon magnetic moment ($M_{ph}$[Niu2021_phononMag, Xue2025_Extrinsic, CeCl3_2022, Bonini2023_CrI3_TRS]. As detailed in Ref. [Niu2021_phononMag], $M_{ph,z}$ comprises two contributions: one from the Born effective charges and the other from the Berry curvature of electrons in both nuclear coordinates (MBC) and momentum space ($\Omega_{e}$), i.e.,
$M_{ph,z}=\frac{e}{2M_{\kappa}}l_{z,\nu\bm{q}}\int d\bm{k}\left\{\Omega_{k_{x}u_{y}}-\Omega_{k_{y}u_{x}}+\Omega_{e,k_{x}k_{y}}G_{u_{x}u_{y}}\right\}.$ (4)
@@
The first two terms collectively integrate to the Born effective charge, whereas the third term is equivalent to the product of the conventional electronic and molecular Berry curvatures. In css, the metallic nature and the preservation of $\mathcal{P}$ result in a vanishing Born effective charge. Given that $G$ is independent of $\bm{k}$ and the integral $\int d\bm{k}\Omega_{e,k_{x}k_{y}}$ is proportional to the anomalous Hall conductivity $\sigma^{\text{A}}_{xy}$, the expression for the phonon magnetic moment simplifies to $M_{ph}\propto l_{z,\nu\bm{q}}\cdot G\cdot\sigma^{\text{A}}_{xy}$. Consequently, the large intrinsic Hall conductivity of css (approximately 505 $\Omega^{-1}$ cm-1 [Wang2018_NC]) indicates a strong propensity for intrinsic phonon magnetism in this material. As calculated and shown in Fig. 3, the pronounced difference in the MBC between two modes indicates a larger magnetic moment for the $E_{g}$ phonon.
III Conclusion and Discussion
@@
In summary, we establish an ab initio MBC framework that captures electronic-order–induced symmetry breaking and phonon magnetism in magnetic materials, exemplified by Co3Sn2S2. Remarkably, the energy scale of circularly polarized phonons coincides with that of chiral Weyl fermions, unveiling opportunities for chiral boson–fermion coupling and enabling control of topological states through phononic and magnetic interactions [hernandez2023observation]. This framework offers an efficient and accurate route to modeling magnetic lattice dynamics, readily extendable to full Brillouin-zone calculations and the exploration of emergent collective phenomena. From these results, we derive design principles for engineering large phonon magnetic moments and sizable phonon splitting, which require strong electron–phonon coupling, pronounced SOC, and phonon modes with finite angular momentum. Magnetic materials with flat electronic bands, such as Co3Sn2S2, thus provide an ideal platform for realizing giant phonon splitting and pronounced phonon magnetism.
Beyond Co3Sn2S2, our algorithm also predicts ferromagnetic materials such as FeCo and Co2MnSi, both with $O_{h}$ symmetry, to exhibit FM-order–induced phonon splitting and phonon magnetism in their lattice dynamics. The MBC-induced maximum splitting of the $E$ mode along the $C_{4}$-invariant path is calculated to be 1.86 cm-1 for FeCo and 0.16 cm-1 for Co2MnSi. Further details are provided in the Supplementary Materials. We note that larger splittings may be observed experimentally due to additional effects such as Fano resonances or electronic correlations.
IV Methods
The electronic structure and the density functional perturbation theory (DFPT) calculation is performed by the Quantum Espresso package [QE1, QE2, QE3]. The norm-conserving pseudo-potential of the Perdew-Burke-Ernzerhof (PBE) exchange–correlation functional [perdew_generalized_1996, dojo, ONCV_Hamann] was used. The kinetic energy cutoff is set to 90 Ry. The Brillouin zone was sampled using a $8\times 8\times 8$ Monkhorst–Pack mesh. The MBC is calculated under the first-order-perturbation framework and calculated by the Wannier-based electron-phonon coupling vortex with modified EPW code [mostofi2008wannier90, mostofi2014updated, EPW1, EPW2, EPW3]. The MBC is Wannier-interpolated on a $80\times 80\times 80$ fine $k$-mesh. The core code will be published along with the manuscript.
Acknowledgments
@@
We acknowledge the helpful discussion with Haijun Zhang, Fuyi Wang, Xi Dai, Yang Gao, Yafei Ren, Shang Ren, Run Yang and Martin Dressel.
Foundings:
We acknowledge the support from National Key R&D Project (Grant Nos. 2023YFA1407400 and 2024YFA1409200), and the National Natural Science Foundation of China Grant No. 12374165.
Author contributions:
T.Z. conceived of the presented idea. T.Z. and S.Z. developed the theory and performed the first-principles calculations. All authors discussed the results and contributed on the manuscript.
Competing interests:
The authors declare that they have no competing interests.
Data availability:
All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials. The code is available from the corresponding author upon reasonable request.